OpenStudy (anonymous):

Find the product of (x - 3)^2.

OpenStudy (anonymous):

@ObeySoph▲

OpenStudy (zzr0ck3r):

\((a+b)^2=a^2+2ab+b^2\)

OpenStudy (texaschic101):

(x - 3)^2 (x - 3)(x - 3) Use the FOIL method (first,outside,inside,last) first -- multiply first terms in each set....x * x = x^2 outside -- multiply outside terms in each set...x * -3 = -3x inside -- multiply inside terms in each set..-3 * x = -3x last -- multiply last terms in each set...-3 * -3 = 9 now put it all together.. x^2 - 3x - 3x + 9 -- combine like terms x^2 - 6x + 9

OpenStudy (zzr0ck3r):

what part of texas?

OpenStudy (texaschic101):

east

OpenStudy (zzr0ck3r):

born in tyler

OpenStudy (texaschic101):

about 200 miles from Houston

OpenStudy (zzr0ck3r):

close...

OpenStudy (texaschic101):

Tyler is a little over an hr from me

OpenStudy (anonymous):

thanks texas

OpenStudy (zzr0ck3r):

np

OpenStudy (zzr0ck3r):

oh her

OpenStudy (zzr0ck3r):

jk:P

OpenStudy (texaschic101):

sure thing style :)

OpenStudy (texaschic101):

lol....you crack me up zzr0ck3r

OpenStudy (zzr0ck3r):

:)

OpenStudy (texaschic101):

:)